OpenStudy (anonymous):

The polynomial given below has______ roots(s). 4x^2+8x+5 A. one negative and one positive B. two negative C. two positive D. two complex

OpenStudy (campbell_st):

use the discriminant to determine the nature of the roots \[\Delta = b^2 - 4ac\] in your question a = 4, b = 8 and c = 5 the conditions discriminant > 0 2 real discriminant = 0 a repeated root discriminant < 0 2 complex roots this should help

OpenStudy (anonymous):

i got 16

OpenStudy (campbell_st):

is it positive 16 or negative 16..?

OpenStudy (anonymous):

+

OpenStudy (campbell_st):

I think you need to check it...

OpenStudy (campbell_st):

\[8^2 - 4 \times 4 \times 5 = \]

OpenStudy (anonymous):

snap -16

OpenStudy (campbell_st):

ok... so go and look at the conditions for the discriminant.... and see what the answer means...

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

because the 16 is -

OpenStudy (campbell_st):

nope.... go and read my post above about the values for the discriminant....

OpenStudy (anonymous):

complex root

OpenStudy (campbell_st):

thats correct...